OpenStudy (anonymous):

The ______ of a triangle is a segment that connects the midpoints of two sides of the triangle

OpenStudy (anonymous):

mid-segment ? O.o

OpenStudy (anonymous):

Ty allot!!

OpenStudy (anonymous):

I am very bad at math

OpenStudy (anonymous):

we all are

OpenStudy (anonymous):

one more??

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

A point is ____ from two objects if it is the same distance from the objects

OpenStudy (anonymous):

equidistant

OpenStudy (anonymous):

well your probably tired of these questions! so ill stop asking u if u want

OpenStudy (anonymous):

u really did help me though which i really do appriciate

OpenStudy (anonymous):

these answers should be in your book's theory,or in definitions but i will help

OpenStudy (anonymous):

i am doing online school which i really do hate i would rather go to regular school and be tought from person to person not from me looking at a computer screen

OpenStudy (anonymous):

thats awful i dont like cyber schools

OpenStudy (anonymous):

yeah it sucks... soo u said ur up for one more so one second

OpenStudy (anonymous):

yeah okay

OpenStudy (anonymous):

the distance from a point to a line is the length of the _____ segment from the point tothe line

OpenStudy (anonymous):

i am confused here @ganeshie8 would that be perpendicular? i see many answers

ganeshie8 (ganeshie8):

yes perpendicular fits perfectly :)

OpenStudy (anonymous):

Awesome!!! ^_^

OpenStudy (anonymous):

when three or more lines intersect at one point, they are _____

OpenStudy (anonymous):

ima try and figure this out but try and help but dont give me the answer

OpenStudy (anonymous):

okay then

OpenStudy (anonymous):

ok lets do this!

ganeshie8 (ganeshie8):

http://en.wikipedia.org/wiki/Concurrent_lines

OpenStudy (anonymous):

I cant think of anything my mind is at blank

OpenStudy (anonymous):



OpenStudy (anonymous):

haha

OpenStudy (anonymous):

ok so i think they are concurrent

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

IT WAS ALL ME

OpenStudy (anonymous):

gane definently didnt help....

OpenStudy (anonymous):

one more bro

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

DF bisects m<EDG. find FG

OpenStudy (anonymous):

is there a image? or value?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

i helped u by not helping @HelpPlease07 :P

OpenStudy (anonymous):

i was kidding!! :)

ganeshie8 (ganeshie8):

since DF is a bisector, every point on DF will be at same distance to both the arms : FE = FG n+5 = 2n-6 solve n first

OpenStudy (anonymous):

Ummm

OpenStudy (anonymous):

Im sorry i cant do this

OpenStudy (anonymous):

... how old are you @HelpPlease07 i hate cyberschools @ganeshie8

OpenStudy (anonymous):

Im 16

OpenStudy (anonymous):

Conexus

OpenStudy (anonymous):

okay so you are now solving for n n+5 = 2n-6 so put n to right side and bring 6 to left side

OpenStudy (anonymous):

so you get 5+6 = 2n -n because when you switch the sides,the sign is changed so n=11

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

so 6+5 = n

OpenStudy (anonymous):

= 11

OpenStudy (anonymous):

= N

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

HAHAHA IM DOING MATH!!!

OpenStudy (anonymous):

so now you have to find FG and FG = 2n-6 so put value of n into 2n-6 and tell me the final answer

OpenStudy (anonymous):

6???

OpenStudy (anonymous):

OH ITS 11

OpenStudy (anonymous):

sorry my bad i get it confused with 6

OpenStudy (anonymous):

22 - 6 = 16???

OpenStudy (anonymous):

yes good

OpenStudy (anonymous):

you should become a teacher

OpenStudy (anonymous):

naah students are pain in the.. you know :P

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Well awesome dude you helped me allot!!!

OpenStudy (anonymous):

its 4:00 AM right now so ima hit the hay

OpenStudy (anonymous):

lol sure night night